OpenStudy (anonymous):

Which of the following are equivalent to the function y = 4cos x - 2? Check all that apply. A. y = -4cos x + 2 B. y = 4sin(x +π/2 ) - 2 C. y = 4sin(x - π/2) - 2 D. y = 4cos(-x) - 2

myininaya (myininaya):

cos(x)=cos(-x) so D

OpenStudy (anonymous):

just d?

myininaya (myininaya):

i didnt bother going through the rest but you can and i will check them for you

myininaya (myininaya):

here i will write down some clues for you \[\sin(\frac{\pi}{2}-\theta)=\cos(\theta)\] \[\cos(\frac{\pi}{2}-\theta)=\sin(\theta)\] \[\sin(-\theta)=-\sin(\theta)\]

OpenStudy (anonymous):

so b and c as well?

myininaya (myininaya):

cosx=sin(pi/2-x)=sin(-x+pi/2)=sin(-[x-pi/2]) =-sin(x-pi/2) so neither c or b unless c as a negative in front of it